OpenStudy (anonymous):

Choose the correct simplification of the expression (d3)5

OpenStudy (anonymous):

d8 d15 d243 d

jimthompson5910 (jim_thompson5910):

assuming this is (d^3)^5, you would multiply the exponents

jimthompson5910 (jim_thompson5910):

because something like (d^3)^2 = d^(3*2) = d^6

OpenStudy (anonymous):

d^6 isnt an option

jimthompson5910 (jim_thompson5910):

that's just an example

jimthompson5910 (jim_thompson5910):

sorry should have said that

OpenStudy (anonymous):

oh i got d8.. is that right?

jimthompson5910 (jim_thompson5910):

the exponents are 3 and 5, they multiply to ???

jimthompson5910 (jim_thompson5910):

you don't add them, you multiply them

OpenStudy (anonymous):

15

OpenStudy (anonymous):

so its d15

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)4

OpenStudy (anonymous):

its a small four on the top of the 3

jimthompson5910 (jim_thompson5910):

(3x)^4 = (3^4)*x^4 (3x)^4 = 81x^4

OpenStudy (anonymous):

Choose the correct simplification of the expression (x2y)2.

jimthompson5910 (jim_thompson5910):

(x^2)^2 = ???

OpenStudy (anonymous):

4? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^4 actually

jimthompson5910 (jim_thompson5910):

so \[\Large (x^2y)^2 = x^4y\]

OpenStudy (anonymous):

x4y3 x6y6 x4y2 xy4

OpenStudy (anonymous):

those r all the options @jim_thompson5910

jimthompson5910 (jim_thompson5910):

oh sry, i meant to say \[\Large (x^2y)^2 = x^4y^2\]